OpenStudy (anonymous):

3/5 of the 30 days in June were sunny. How many day were sunny? a) 3 b) 5 c) 12 d) 18

OpenStudy (anonymous):

3/5*30=18

OpenStudy (anonymous):

Last Question jenny has 1 liter containers, and each container is 2/3 full of sport drinks. How much sport drinks is in 5 containers?

OpenStudy (anonymous):

2/3*5=3.33 repeted.

OpenStudy (anonymous):

in 1 container there is 3.33 liters. 3.33*5=16.65liters. the awnser is 16.65. I only did it for one container.